OpenStudy (anonymous):

Two times a number, x, plus 3 times a number, y, equals 50. Four times x minus 2 times y equals 4. What are the numbers? x = 7, y = 12 x = 19, y = 4 x = 10, y= 18 x = -11, y = 24

OpenStudy (butterflydreamer):

First, you can create two system of equations... given the information in the question :) So what would these two equations be?

OpenStudy (anonymous):

\[2(x) + 3(y) = 50; 4(x) - 2(y) =4\]

OpenStudy (anonymous):

Solve the first one

OpenStudy (anonymous):

\[3(y) = -2(y) +50\]